# Twists of Superconformal Algebras

### v2

@@ 8.1. Superconformal Prefactorization Algebras
We will follow the conventions of [CG1, CG2], so our prefactorization algebras take values in cochain complexes of differentiable vector spaces. This setting allows us to talk about smooth action of a Lie group on a prefactorization algebra, with a Lie algebra acting by differential operators. The experienced reader, however, will see how the idea of such symmetry could be ported to other settings.
@@
We note the following, which follows by inspection (at least for the reader of sections 3.7 and 4.8 of [CG1] or of section 2 of [ElliottSafronov]).
@@
[…]
where $zU_{1}$, for instance, denotes the translation of $U_{1}$ by $z$. (See definition 3.7.1 of [CG1].) If we take $V$ to be the whole manifold and pick a conformal state (or expected value map) $\langle-\rangle:\mathcal{F}^{\mathcal{Q}}(M)\to\mathbb{C}$, this equivariance implies
@@ 8.2. Algebras of Operator Products
We overview one way to encode the most important operators and their products for a field theory on the manifold $\mathbb{R}^{n}$. The essential idea is that one specifies the operators supported in any disk
@@
A detailed treatment of this notion appears in section 4.8 of [CG1] and in section 2 of [ElliottSafronov].
@@
A technical issue is to formulate in what sense $m$ is a smooth function, which requires finding a setting where the codomain $\mathrm{Hom}$ has the necessary structure. In [CG1] it is shown that differentiable vector spaces provide a sufficient setting and it is shown that most topological vector spaces are differentiable vector spaces.
When a Lie group $G$ acts smoothly by isometries on $\mathbb{R}^{n}$, it induces an action on the configuration spaces $\mathrm{Disk}_{r_{1},\ldots,r_{k};R}$, so one can talk about a $n$-disks algebra $\mathcal{A}$ being $G$-equivariant. Similarly, one can formulate how a dg Lie algebra can act by derivations on $\mathcal{A}$.
@@ 8.3. Obtaining Vertex Algebras or $\mathbb{E}_{k}$-Algebras
Under some reasonable hypotheses, one can obtain a vertex algebra or an $\mathbb{E}_{k}$ algebra (i.e., algebra over the little $k$-disks operad) from the observables of a field theory, whether encoded as a prefactorization algebra or as a disk-algebra. Precise results appear in Chapter 5 of [CG1] and Section 2 of [ElliottSafronov]. Here we will state their consequences.
@@ 8.3.1. Obtaining $\mathbb{E}_{k}$-Algebras
Let us discuss the circumstances under which the restricted twisted theory $j^{*}\mathcal{F}^{\mathcal{Q}}$ on an affine subspace $\mathbb{R}^{m,n}$ given by the embedding $j$ is topological. Corollary 2.30 of [ElliottSafronov] gives two conditions that guarantee that $j^{*}\mathcal{F}^{\mathcal{Q}}$ determines an algebra over the little $m+n$-disks operad – or more colloquially, that the twist looks like a topological field theory along $\mathbb{R}^{m,n}$. These conditions are
@@
Corollary 2.39 applies to the case that $n=0$. It says that one gets a framed $\mathbb{E}_{m}$-algebra if, furthermore, the subalgebra $\mathfrak{so}(m)$ is $\mathcal{Q}$-exact. Such a situation appears not infrequently in Euclidean signature. (An analog of this result should hold for mixed signature as well, using the techniques of [ElliottSafronov].)
Let us make this application of the results of [ElliottSafronov] more precise.
@@
  • the group of translations of $j(\mathbb{R}^{a,b})$ is contained in $Z_{\mathcal{Q}}$.
@@
  • $[\mathcal{Q},-]$ defines a linear isomorphism $\mathfrak{h}^{\prime}\to\mathfrak{h}$.
  • $[\mathfrak{h}^{\prime},\mathfrak{h}^{\prime}]=0$.
  • $[\mathfrak{h},\mathfrak{h}^{\prime}]\subseteq\mathfrak{h}^{\prime}$, and as an $\mathfrak{h}$-representation $\mathfrak{h}^{\prime}$ is isomorphic to the adjoint representation.
@@
Now, Corollary 2.30 and Theorem 2.38 of [ElliottSafronov] can be applied to the prefactorization algebra $j^{*}\mathrm{Obs}^{\mathcal{Q}}$ in the presence of appropriate potentials. We obtain the following as a direct application of these results.
@@
Let $j\colon\mathbb{R}^{a,b}\hookrightarrow C(\mathbb{R}^{p,q})$ be a $\mathcal{Q}$-compatible subspace, and suppose that the group of translations of $j(\mathbb{R}^{a,b})$ is contained in $B_{\mathcal{Q}}\subseteq Z_{\mathcal{Q}}$. Let $G\subseteq B_{\mathcal{Q}}$ be a subgroup acting on the image of $j$. Let $\mathfrak{g}\subseteq\mathfrak{b}^{\mathbb{C}}_{\mathcal{Q}}$ denote the complexified Lie algebra of $G$, and let $\mathfrak{g}^{\prime}$ be a $\mathcal{Q}$-potential for $\mathfrak{g}$. Finally, suppose that the factorization map $j^{*}\mathrm{Obs}^{\mathcal{Q}}(B_{r}(0))\to j^{*}\mathrm{Obs}^{\mathcal{Q}}(B_{R}(0))$ associated to the inclusion of concentric balls is a quasi-isomorphism for all $0<r<R$. Then $j^{*}\mathrm{Obs}^{\mathcal{Q}}(B_{1}(0))$ canonically carries the structure of a $G\ltimes\mathbb{E}_{a+b}$-algebra.
Corollary 8.13 (c.f. [ElliottSafronov, Corollary 3.40]).
If there is an embedding ×$\mathbb{R}^{\times}\hookrightarrow G$ whose image acts on the image of $j$ by dilations, then the condition that $j^{*}\mathrm{Obs}^{\mathcal{Q}}(B_{r}(0))\to j^{*}\mathrm{Obs}^{\mathcal{Q}}(B_{R}(0))$ is a quasi-isomorphism holds automatically.
@@ 8.3.2. Obtaining Vertex Algebras
    The group $\mathrm{ISO}(j(\mathbb{R}^{2}))$ of isometries of the image of $j$ is contained in $Z_{\mathcal{Q}}$.
@@
This will be a direct consequence of [CG1, Theorem 2.2.1], and we refer the reader to that reference for a discussion of the meaning of the technical conditions that appear in the statement of this theorem. In particular, the “tameness” condition that we have not spelled out here is condition (iii) in loc. cit.
@@ 8.4. The Approach of Saberi–Williams to Schur twists
Saberi and Williams [SaberiWilliams] take a different approach to understanding the occurrence of vertex algebras from Schur twists of 4d $\mathcal{N}=2$ theories (in Euclidean signature), an approach that is fundamentally constructive and hence produces a number of explicit examples. They focus on factorization algebras built with the methods of [CG1, CG2], and so their examples arise as the observables of perturbative supersymmetric theories or as current algebras (i.e., local symmetries) of such theories. Here we will review their construction and mention some key results, and then we will explain how their rich physical examples offer examples of the framework of this paper.
@@
This local Lie algebra $\mathfrak{X}_{2}$ generates a current algebra on $\mathbb{C}^{2}$ that is an interesting two-dimensional analog of the Virasoro algebra; there are local cocycles analogous to the central extension (or Schwinger term) that determines the Virasoro algebra. This current algebra is an example of an enveloping factorization algebra, and it is a central example for their work. (See Section 5.4 of [SaberiWilliams] for details.) This proposition thus ensures that a twisted 4d $\mathcal{N}=2$ gauge theory (even after quantization) has an enormous symmetry algebra via a two-dimensional version of the Virasoro algebra.
@@
Their techniques apply to essentially all the Lagrangian $\mathcal{N}=2$ supersymmetric field theories. Throughout [SaberiWilliams] there are detailed computations that match the results of Beem et al.
We now point out that their examples match the hypotheses of Theorem 8.15, where $j$ embeds $\mathbb{C}$ into $\mathbb{C}^{2}$ by $j(z)=(z,0)$: the vertex algebras they produce by “localization” are examples of our abstract result. In particular, we see that many examples of Beem et al fit into our framework, thanks to [SaberiWilliams].
[…]

### v3

@@ 8.1. Superconformal Prefactorization Algebras
We will follow the conventions of [15, 16], so our prefactorization algebras take values in cochain complexes of differentiable vector spaces. This setting allows us to talk about smooth action of a Lie group on a prefactorization algebra, with a Lie algebra acting by differential operators. The experienced reader, however, will see how the idea of such symmetry could be ported to other settings.
@@
We note the following, which follows by inspection (at least for the reader of sections 3.7 and 4.8 of [15] or of section 2 of [23]).
@@
$\mathcal{F}^{\mathcal{Q}}(U_{1})\otimes\mathcal{F}^{\mathcal{Q}}(U_{2})$$\mathcal{F}^{\mathcal{Q}}(V)$$\mathcal{F}^{\mathcal{Q}}(zU_{1})\otimes\mathcal{F}^{\mathcal{Q}}(zU_{2})$$\mathcal{F}^{\mathcal{Q}}(zV)$$\scriptstyle{\cong}\scriptstyle{\cong}$
where $zU_{1}$, for instance, denotes the translation of $U_{1}$ by $z$. (See definition 3.7.1 of [15].) If we take $V$ to be the whole manifold and pick a conformal state (or expected value map) $\langle-\rangle:\mathcal{F}^{\mathcal{Q}}(M)\to\mathbb{C}$, this equivariance implies
@@ 8.2. Algebras of Operator Products
We overview one way to encode the most important operators and their products for a field theory whose underlying manifold is $\mathbb{R}^{n}$. The essential idea is that one specifies the operators supported in any disk
@@
A detailed treatment of this notion appears in section 4.8 of [15] and in section 2 of [23].
@@
A technical issue is to formulate in what sense the multiplication $m$ is a smooth function, which requires finding a setting where the codomain $\mathrm{Hom}$ has the necessary structure. In [15] it is shown that differentiable vector spaces provide a sufficient setting and it is shown that most topological vector spaces are differentiable vector spaces.
A dg Lie algebra can act by derivations on a $n$-disks algebra $\mathcal{A}$. Here is the key example for this paper.
@@ 8.3. Obtaining Vertex Algebras or $\mathbb{E}_{k}$-Algebras
Under some reasonable hypotheses, one can obtain a vertex algebra or an $\mathbb{E}_{k}$ algebra (i.e., algebra over the little $k$-disks operad) from the observables of a field theory, whether encoded as a prefactorization algebra or as a disk-algebra. Precise results appear in Chapter 5 of [15] and Section 2 of [23]. Here we will state their consequences.
@@ 8.3.1. Obtaining $\mathbb{E}_{k}$-Algebras
Let us discuss the circumstances under which the restricted twisted theory $j^{*}\mathcal{F}^{\mathcal{Q}}$ on an affine subspace $\mathbb{R}^{m,n}$ given by the embedding $j$ is topological. Corollary 2.30 of [23] gives two conditions that guarantee that $j^{*}\mathcal{F}^{\mathcal{Q}}$ determines an algebra over the little $m+n$-disks operad – or more colloquially, that the twist looks like a topological field theory along $\mathbb{R}^{m,n}$. These conditions are
@@
Corollary 2.39 applies to the case that $n=0$. It says that one gets a framed $\mathbb{E}_{m}$-algebra if, furthermore, the subalgebra $\mathfrak{so}(m)$ is $\mathcal{Q}$-exact. Such a situation appears not infrequently in Euclidean signature. (An analog of this result should hold for mixed signature as well, using the techniques of [23].)
Let us make this application of the results of [23] more precise.
@@
  • There is a subgroup $T_{\mathcal{Q}}$ of $Z_{\mathcal{Q}}$ isomorphic to $\mathbb{R}^{a+b}$ whose action on $j(\mathbb{R}^{a,b})$ corresponds to the translation on $\mathbb{R}^{a,b}$.
@@
  • $\mathcal{Q},-]$ defines a linear isomorphism $\mathfrak{h}^{\prime}\to\mathfrak{h}$.
  • $\mathfrak{h}^{\prime},\mathfrak{h}^{\prime}]=0$.
  • $\mathfrak{h},\mathfrak{h}^{\prime}]\subseteq\mathfrak{h}^{\prime}$, and as an $\mathfrak{h}$-representation $\mathfrak{h}^{\prime}$ is isomorphic to the adjoint representation.
@@
Now, Corollary 2.30 and Theorem 2.38 of [23] can be applied to the prefactorization algebra $j^{*}\mathrm{Obs}^{\mathcal{Q}}$ in the presence of appropriate potentials. We obtain the following as a direct application of these results.
@@
Let $j\colon\mathbb{R}^{a,b}\hookrightarrow C(\mathbb{R}^{p,q})$ be a $\mathcal{Q}$-compatible subspace, and suppose that the group $T_{\mathcal{Q}}$ from Definition 8.9 of translations of $j(\mathbb{R}^{a,b})$ is contained in $B_{\mathcal{Q}}\subseteq Z_{\mathcal{Q}}$. Let $G\subseteq B_{\mathcal{Q}}$ be a subgroup acting on the image of $j$. Let $\mathfrak{g}\subseteq\mathfrak{b}^{\mathbb{C}}_{\mathcal{Q}}$ denote the complexified Lie algebra of $G$, and let $\mathfrak{g}^{\prime}$ be a $\mathcal{Q}$-potential for $\mathfrak{g}$. Finally, suppose that the factorization map $j^{*}\mathrm{Obs}^{\mathcal{Q}}(B_{r}(0))\to j^{*}\mathrm{Obs}^{\mathcal{Q}}(B_{R}(0))$ associated to the inclusion of concentric balls is a quasi-isomorphism for all $0<r<R$. Then $j^{*}\mathrm{Obs}^{\mathcal{Q}}(B_{1}(0))$ canonically carries the structure of a $G\ltimes\mathbb{E}_{a+b}$-algebra.
Corollary 8.13 (c.f. [23, Corollary 3.40]).
If there is an embedding $\mathbb{R}^{\times}\hookrightarrow G$ whose image acts on the image of $j$ by dilations, then the condition that $j^{*}\mathrm{Obs}^{\mathcal{Q}}(B_{r}(0))\to j^{*}\mathrm{Obs}^{\mathcal{Q}}(B_{R}(0))$ is a quasi-isomorphism holds automatically.
@@ 8.3.2. Obtaining Vertex Algebras
    The group $Z_{\mathcal{Q}}$ contains a subgroup isomorphic to $\mathrm{ISO}(j(\mathbb{R}^{2}))$ acting by isometries on the image of $j$.
@@
This result is be a direct consequence of [15, Theorem 2.2.1], and we refer the reader to that reference for a discussion of the meaning of the technical conditions that appear in the statement of this theorem. In particular, the “tameness” condition that we have not spelled out here is condition (iii) in loc. cit.
@@ 8.4. The Approach of Saberi–Williams to Schur twists
Saberi and Williams [40] take a different approach to understanding the occurrence of vertex algebras from Schur twists of 4d $\mathcal{N}=2$ theories (in Euclidean signature), an approach that is fundamentally constructive and hence produces a number of explicit examples. They focus on factorization algebras built with the methods of [15, 16], and so their examples arise as the observables of perturbative supersymmetric theories or as current algebras (i.e., local symmetries) of such theories. Here we will review their construction and mention some key results, and then we will explain how their rich physical examples offer examples of the framework of this paper.
@@
This local Lie algebra $\mathfrak{X}_{2}$ generates a current algebra on $\mathbb{C}^{2}$ that is a kind of two-dimensional analog of the Virasoro algebra; there are local cocycles analogous to the central extension (or Schwinger term) that determines the Virasoro algebra. (Perhaps some readers may prefer to view it as a two-dimensional analog of the $N=2$ superconformal algebra, rather than the Virasoro algebra itself, since it involves a derived direction.) This current algebra is an example of an enveloping factorization algebra, and it is a central example for their work. (See Section 5.4 of [40] for details.) This proposition thus ensures that a twisted 4d $\mathcal{N}=2$ gauge theory (even after quantization) has an enormous symmetry algebra via a two-dimensional version of the Virasoro algebra.
@@
Their techniques apply to essentially all the Lagrangian $\mathcal{N}=2$ supersymmetric field theories. Throughout [40] there are detailed computations that match the results of Beem et al.
We now point out that their examples match the hypotheses of Theorem 8.15, where $j$ embeds $\mathbb{C}$ into $\mathbb{C}^{2}$ by $j(z)=(z,0)$: the vertex algebras they produce by “localization” are examples of our abstract result. In particular, we see that many examples of Beem et al fit into our framework, thanks to [40].